Arturo started to deposit 2,000 monthly in a fund that pays 8% compounded quarterly how much will be in the fan after 20 years

1
Expert's answer
2022-03-29T17:00:07-0400

"FV=pmt[\frac{(1+\frac{r}{m})^{nm}-1}{\frac{r}{m}}]"

Where FV is the future value

Pmt is the monthly deposits, 2000

R is the rate, 8%

N is the number of years, 20

M is the number of times compounded, 4


"\therefore FV=2000 [\frac{(1+\frac{8\%}{4})^{4×20}-1}{\frac{8\%}{4}}]"

"FV=387543.92"